# Bibliography management: natbib package

### v1

@@ A.2 Iterative smoothing
It is also necessary to choose values for $\Delta$ and $N$ to be used in the iterative smoothing of each dataset. To reiterate, the aim is to obtain a smooth function which has a better fit to a particular dataset, without fitting the small-scale noise.
@@
In this case, since we are only interested in mutual consistency and not in reconstruction per se, we have a little more freedom in the choice $\Delta$181818as explained in Section 3.2, the aim of the iteratively-smoothed function is to pick up some of the features of the particular dataset and to be a better fit than the concordance cosmology. The effects from using iterative smoothing to generate a template function will be moderated by the Crossing Statistics that are subsequently applied: If there are features inherited by the iterative smoothing template function that are due only to the small-scale noise characteristics of the particular dataset used, and not a common underlying model, then the subsequent versions with smooth deformations from the applied crossing functions will not be able to reconcile the smoothed function as a good fit to the other datasets, without much higher order in crossing functions.. Thus, after testing various values of smoothing length and the number of iterations required for convergence, in the case of Pantheon+ we adopt a value of $\Delta=0.4$, slightly larger than in the literature, to ensure convergence within a more reasonable number of iterations. The DES-SN5YR dataset has not only slightly more supernovae, but also a much smaller redshift range than Pantheon+ (about half), and so the density of datapoints is higher. For this reason, we opt for a slightly smaller value of $\Delta=0.3$. In the initial analysis, we do not use the iteratively-smoothed function from DES-SN5YR as a template function for other datasets, because of the limited redshift range, but we do perform a cross-check using the dataset itself. The Union3 dataset has much fewer data points than either of the other two datasets, and so the value of $\Delta$ must be increased in order that each point of the smoothing function be sensitive to at least one datapoint. We use $\Delta=0.6$.
@@
In addition to these choices of hyperparameters, we discovered that there were some convergence issues when using iterative smoothing on the Union3 dataset. This was found to be due to the relatively low number of binned datapoints, and a relatively strongly negative off-diagonal element in the inverse of the covariance matrix for the correlation between the $3^{rd}$ and $4^{th}$ bins. The element itself is not the most strongly negative one in the inverse covariance matrix, but the way that the iterative smoothing method operates means that it causes a problem. In the $(k+1)^{th}$ iteration, the previous estimate has the following term added to it:
@@
where the numerator is a measure of the residuals present in the previous iteration’s estimate, each weighted by the distribution function $W_{i}(z)$ and a measure of their uncertainty (or freedom given by the data) in the form of the inverse covariance matrix. The denominator is a normalisation factor, meant to ensure that at each step the contribution from this factor is of the order of the residuals, to ensure convergence. It comprises the weighting function $W_{i}(z)$ and $\mathbb{1}^{T}\cdot{\rm C}^{-1}$, which is a vector whose elements are each the sum of one column of the inverse covariance matrix.
@@ Appendix B Contour plot summary
To avoid the unwieldy proliferation of contour plots, in this appendix we summarise the important information for each part of the analysis. For the sake of space, we abbreviate the dataset to which the deformed functions are fit as “P+”, “U3”, “D5” for Pantheon+, Union3 and DES-SN5YR respectively, as well as ‘template function’ as “T.F.”, in the tables that follow. In the case of the initial analysis, including the data over the full range of redshift, we also report the contours plots for the $3^{rd}$ and $4^{th}$ order deformations of the Union3 and DES-SN5YR template functions in the relevant sections (see Figs 1314 and 15).
@@ B.1.1 $\Lambda$CDM best fit template function
Table 8: Number of $\sigma$ at which all the crossing hyperparameters at a particular order are consistent with the standard values, for each Flat $\Lambda$CDM best fit template function, fit to each dataset.
B.1.2 Iterative smoothing template function
Iterative Smoothing T.F.: Pantheon+ Union3 DES-SN5YR
@@ B.1.2 Iterative smoothing template function
Table 9: Number of $\sigma$ at which all the crossing hyperparameters at a particular order are consistent with the standard values, for both iterative smoothing template functions, fit to each dataset.
Figure 13: Contour plots for the hyperparameters of the Union3 $\Lambda$CDM best fit template function, deformed to fit to all datasets. The contours from fits to Pantheon+, Union3 and DES-SN5YR data are shown in blue, pink and orange, respectively. We show the results at maximum crossing function order of $3$ (upper panel) and $4$ (lower panel).
Figure 14: Contour plots for the hyperparameters of the DES-SN5YR $\Lambda$CDM best fit template function, deformed to fit to all datasets. The contours from fits to Pantheon+, Union3 and DES-SN5YR data are shown in blue, pink and orange, respectively. We show the results at maximum crossing function order of $3$ (upper panel) and $4$ (lower panel).
Figure 15: Contour plots for the hyperparameters of the Union3 iterative smoothing template function, deformed to fit to all datasets. The contours from fits to Pantheon+, Union3 and DES-SN5YR data are shown in blue, pink and orange, respectively. We show the results at maximum crossing function order of $3$ (upper panel) and $4$ (lower panel).
@@ B.2.1 $\Lambda$CDM best fit template function
Table 10: Number of $\sigma$ at which all the crossing hyperparameters at a particular order are consistent with the standard values, for each Flat $\Lambda$CDM best fit template function, fit to each dataset with the low $z_{max}$ cut-off.
B.2.2 Iterative smoothing template function
@@
Table 11: Number of $\sigma$ at which all the crossing hyperparameters at a particular order are consistent with the standard values, for each iterative smoothing template function, fit to each dataset with the low $z_{max}$ cut-off.

### v2

@@ A.2 Iterative smoothing
It is also necessary to choose values for $\Delta$ and $N$ to be used in the iterative smoothing of each dataset. To reiterate, we aim to obtain a smooth function which has a viable fit to the dataset, without fitting the small-scale noise and independent of cosmological parameters.
@@
In this case, since we are only interested in mutual consistency and not in reconstruction per se, we have a little more freedom in the choice of $\Delta$ [^20] As explained in Section 3.2, the aim of the iteratively-smoothed function is to pick up some of the features of the particular dataset and to be a viable fit without using the concordance cosmology parameters. The effects from using iterative smoothing to generate a template function will be moderated by the Crossing Statistics that are subsequently applied: If there are features inherited by the iterative smoothing template function that are due only to the small-scale noise characteristics of the particular dataset used, and not a common underlying model, then the subsequent versions with smooth deformations from the applied crossing functions will not be able to reconcile the smoothed function as a good fit to the other datasets, without much higher order in crossing functions.. Thus, after testing various values of smoothing length and the number of iterations required for convergence, in the case of Pantheon+ we adopt a value of $\Delta=0.4$, slightly larger than in the literature, to ensure convergence within a more reasonable number of iterations. The DES-SN5YR dataset has not only slightly more supernovae, but also a much smaller redshift range than Pantheon+ (about half), and so the density of data points is higher. For this reason, we opt for a slightly smaller value of $\Delta=0.3$. In the initial analysis, we do not use the iteratively-smoothed function from DES-SN5YR as a template function for other datasets, because of the limited redshift range, but we do perform a cross-check using the dataset itself. The Union3 dataset has fewer data points, and so the value of $\Delta$ must be increased in order that each point of the smoothing function be sensitive to at least one datapoint. We use $\Delta=0.6$.
@@
In addition to these choices of hyperparameters, we discovered some convergence issues when using iterative smoothing on the Union3 dataset. This was found to be due to the relatively low number of splined data points, and a relatively strongly negative off-diagonal element in the inverse of the covariance matrix for the correlation between the $3^{rd}$ and $4^{th}$ spline nodes. The element itself is not the most strongly negative one in the inverse covariance matrix, but the way that the iterative smoothing method operates means that it causes a problem. In the $(k+1)^{th}$ iteration, the previous estimate has the following term added to it:
@@
where the numerator is a measure of the residuals present in the previous iteration’s estimate, each weighted by the distribution function $W_{i}(z)$ and the inverse covariance matrix (a measure of the uncertainty/freedom present in the data). The denominator is a normalisation factor, meant to ensure that at each step the contribution from this factor is of the order of the residuals, for convergence. It comprises the weighting function $W_{i}(z)$ and $\mathbb{1}^{T}\cdot{\rm C}^{-1}$, which is a vector whose elements are each the sum of one column of the inverse covariance matrix.
@@ Appendix B Contour plot summary
To avoid the unwieldy proliferation of contour plots, in this appendix we summarise the important information for each part of the analysis. For the sake of space, we abbreviate the dataset to which the deformed functions are fit as “P+”, “U3”, “D5” for Pantheon+, Union3 and DES-SN5YR respectively, as well as ‘template function’ as “T.F.”, in the tables that follow. In the case of the initial analysis, including the data over the full range of redshift, we also report the contours plots for the $3^{rd}$ and $4^{th}$ order deformations of the Union3 and DES-SN5YR template functions in the relevant sections (see Figs 1415 and 16).
@@ B.1 Full redshift ranges
Table 8: $\sigma$-level at which all the crossing hyperparameters at each order are compatible with standard values, for each Flat $\Lambda$CDM best fit template function, fit to each dataset.
Iterative Smoothing: Pantheon+ Union3 DES-SN5YR
@@
Table 9: $\sigma$-level at which all the crossing hyperparameters at each order are compatible with standard values, for both iterative smoothing template functions, fit to each dataset.
Figure 14: Contour plots for the hyperparameters of the Union3 $\Lambda$CDM best fit template function, deformed to fit to all datasets. The contours from fits to Pantheon+, Union3 and DES-SN5YR data are shown in blue, pink and orange, respectively. We show the results at maximum crossing function order of $3$ (upper panel) and $4$ (lower panel), for the full redshift range.
Figure 15: Contour plots for the hyperparameters of the DES-SN5YR $\Lambda$CDM best fit template function, deformed to fit to all datasets. The contours from fits to Pantheon+, Union3 and DES-SN5YR data are shown in blue, pink and orange, respectively. We show the results at maximum crossing function order of $3$ (upper panel) and $4$ (lower panel), for the full redshift range.
Figure 16: Contour plots for the hyperparameters of the Union3 iterative smoothing template function, deformed to fit to all datasets. The contours from fits to Pantheon+, Union3 and DES-SN5YR data are shown in blue, pink and orange, respectively. We show the results at maximum crossing function order of $3$ (upper panel) and $4$ (lower panel), for the full redshift range.
@@ B.2 Low $z_{max}$ cut
Table 10: Number of $\sigma$ at which all the crossing hyperparameters at a particular order are compatible with the standard values, for each Flat $\Lambda$CDM best fit template function, fit to each dataset with the low $z_{max}$ cut-off.
@@
Table 11: Number of $\sigma$ at which all the crossing hyperparameters at a particular order are compatible with the standard values, for each iterative smoothing template function, fit to each dataset with the low $z_{max}$ cut-off.